A famous batsman scored a century in 38 overs.If he could keep up the same run rate,how many runs would be scored in 47.5 overs?
Answer by stanbon(75887)

A famous batsman scored a century in 38 overs.If he could keep up the same run rate,how many runs would be scored in 47.5 covers?
-----
Use a proportion::
r/47.5 = 100/38
----
r = 4750/38
r = 125 runs
----------
Cheers,
Stan H.